1. A can is dropped from rest at a height of 40.0 m above ground. How much time does it spend in the bottom half of its fall?
1
Expert's answer
2022-01-23T16:13:45-0500

Given:

"h=\rm 40\:\rm m"

"g=9.8\:\rm m/s^2"


The time of free falling

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2*40}{9.8}}=2.9\:\rm s"